Determine whether each set of data is linear, quadratic, or neither using finite differences.
X Y
0 -1
1 1
2 3
3 5
4 7

Answer by Edwin McCravy(20054)


X  Y  1st differences  
0 -1 
     1-(-1)=1+1=2
1  1
            3-1=2
2  3
            5-3=2
3  5
            7-5=2
4  7

Since the FIRST differences are all the same -- that is, they 
are all 2's, the data is of a FIRST degree polynomial, which
means it is linear, and follows the linear or FIRST DEGREE 
equation

y = 2x - 1

Edwin